OpenStudy (anonymous):

There are 20 students on the school's student council. A special homecoming dance committee is to be formed by randomly selecting 7 students from student council. How many possible committees can be formed?

OpenStudy (anonymous):

5,040 77,520 390,700,800

OpenStudy (anonymous):

It's permutation -> 390,700,800

OpenStudy (anonymous):

explain

OpenStudy (anonymous):

The committee, so the order is matter!

OpenStudy (anonymous):

@nbouscal help on this one for feed back

OpenStudy (anonymous):

I don't think that order matters. You're just picking a committee. A committee of the same 7 people in different order would be the same committee. So, it's a combination. Combination will be \[ \frac{20!}{7!13!} \]

OpenStudy (anonymous):

ok so solve that and get my answer

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get a

OpenStudy (anonymous):

5040

OpenStudy (anonymous):

\[ \frac{20\times19\times18\times17\times16\times15\times14}{7\times6\times5\times4\times3\times2}=\frac{\not20\times19\times\not18\times17\times16\times15\times\not14}{\not7\times\not6\times\not5\times\not4\times\not3\times\not2}\\ 19\times17\times16\times15\\ 77,520 \]

OpenStudy (anonymous):

ok thank you